# FoMo-Bench: a multi-modal, multi-scale and multi-task Forest Monitoring Benchmark for remote sensing foundation models

### v2

@@ 4 The TalloS dataset
| Split | Number of Classes | Number of Samples |
| Train | 1159 | 161,332 |
| Val | 24 | 4,049 |
| Test | 223 | 25,100 |
| Total | 1160 | 190,481 |
As shown in Fig. 2, the spatial coverage of pre-existing datasets in FoMo-Bench is limited. We improve the diversity and spatial coverage of FoMo-Bench by designing a new dataset, TalloS. This dataset couples manual tree inventories with multispectral and SAR timeseries imagery for a fine grained multi-label classification challenge.
As described in the previous section, various tasks in forest monitoring have required high-resolution data from UAVs but these typically have limited spatial coverage due to human and hardware limits [17, 53, 35, 33]. Satellite-based datasets with fine-grained annotations have previously been lacking for forest monitoring applications.
TalloS is a large-scale multi-modal dataset of satellite-based imagery for tree species multi-label classification. It is based on the global Tallo database [32], containing almost manual georeferenced recordings aggregating information from forest inventories including the tree species as well as tree height, stem diameter, and crown radius. Due to the modest ground sampling distance of publicly accessible satellites (), we design TalloS to concentrate on areas containing a minimum of 50 trees, as regions with lower tree populations lack distinctive signatures in satellite imagery.
TalloS data represent a time series from 2016 to 2023 with a temporal frequency of 15 days. Each timestep consists of a multispectral image from Sentinel-2, a SAR image from Sentinel-1, a digital elevation model (DEM) from Copernicus GLO-30, and meteorological information from ERA5 [27]. Each remote sensing sample of TalloS is mapped to ground-validated and fine grained measurements of trees in its given area, including tree species, genus and family as well as their specific traits. For the purpose of this study, we concentrate on the genus level labels, as distinguishing between species may require higher resolution imagery. Throughout this paper, unless explicitly stated otherwise, we refer to the various genus categories of TalloS as its classes. As shown in Table 1 and Figure 3, TalloS represents a fine-grained classification problem involving more than 1,000 classes with a heavily skewed data distribution. The task is framed as multi-label classification, with the goal of predicting every tree genus that is depicted in the image timeseries.
In Figure 2(b), we illustrate the wide spatial distribution of TalloS as well as the proposed training, validation and test splits. Our goal was to maximize the geographic spatial coverage of each split while ensuring (a) that regions used for training and testing did not overlap, which could lead to inaccurate evaluation, (b) that most classes present in the validation and test sets are also included in the training set. Accommodating both conditions proved challenging due to the limited geographic distribution of many tree genera. The final dataset split is shown in Table 1.
5 FoMo-Net
Building on FoMo-Bench and inspired by the necessity for a generalized foundation model, we introduce FoMo-Net. FoMo-Net is a first attempt to process the most common modalities in the remote sensing domain with a single, sensor-agnostic foundation model, offering a robust foundation model baseline for FoMo-Bench.
Our approach is driven by three core considerations. First, it should maximize flexibility to process various input settings, without relying on specific sets of sensory modalities. Second, it should have the capacity to process information and generate meaningful representations for the whole globe. Third, it should be applicable to a wide range of downstream tasks.
Pre-training data: FoMo-Net pre-training scheme contains rich multi-sensor information, both paired and unpaired, from most parts of the world. In particular, we use the RapidAI4EO dataset as a source of Planet and Sentinel-2 data; TalloS containing DEM, Sentinel-1, and Sentinel-2; SSL4EO-Landsat [58] to acquire global information from Landsat 8 and 9; a combination of all available UAV datasets in FoMo-Bench; and the FiveBillionPixel dataset providing a unique source of Gaofen-2 high resolution satellite imagery.
| Dataset | Sensors | Spatial resolution | Sampling weight |
@@ 5 FoMo-Net
| RapidAI4EO | Planet, Sentinel-2 | Medium, Low | 0.2 |
| TalloS | Sentinel-1, Sentinel-2, DEM | Low | 0.2 |
@@
The datasets used in the FoMo-Net pre-training framework are detailed in Table 2. The proposed framework is able to process any combination of the 36 most common modalities in remote sensing, stemming from Sentinel-1, Sentinel-2, Landsat 8-9, Planet, Gaofen-2 and UAV sensors with ground sampling distance spanning from a few centimeters to per pixel. In this work, a distinct modality is defined as any band with a unique combination of wavelength and spatial resolution. Given its importance in Earth observation, digital elevation models (DEM) are also included in our definition as a specific band. The FoMo-Net pre-training pipeline is illustrated in Figure 4 and will be detailed in the following sections. Additionally, the pseudocode of the pre-training pipeline is summarized in Appendix A.
@@
Approach to variable spectral bands: Let be a set of datasets, and be a set of spectral bands, where where and represent respectively the height and width dimensions. At each iteration, a training batch contains a variable number of spectral bands, each sampled with respective probabilities . In our experiments, we set equal to the frequency with which the -th band occurs in (see Table 2).
@@
Each band is tokenized into patches of size according to the transformation , so that . The tokenized input is then embedded using a linear transformation learnt during the optimization process as . We considered two projection setups: The first (noted FoMo-Net1) uses a single linear projection for all spectral bands, while the second (denoted FoMo-Netm) projects each spectral band with its own linear projection. Our initial experiments suggest that FoMo-Net1 yields significantly better results.
The input to our model, then, is provided as:
@@
Spectral band MAE: The next step of the procedure follows a masked autoencoding (MAE) framework [25] that is shared by various high-performing foundation models for remote sensing [60, 8, 52, 67, 16, 62, 28, 23]. However, our method differs from a typical MAE in two key aspects (see Fig. 4). First, we introduce a random spectral band selection by sampling in both and , representing the datasets and spectral bands, respectively. This approach promotes modeling interactions between a wide range of band combinations, creating a highly flexible encoder able to process any set of spectral bands. Next, we incorporate both spectral and positional embeddings, allowing the model to capture information regarding the spatial position of the patch and its spectral band, as detailed in Eq. 1. Both types of embeddings are trainable and randomly initialized. Given the randomness in the input space, the spectral embedding is of utmost importance as no other information regarding the nature of the input is provided.
@@
Batch gradient accumulation: The combination of the proposed modality sampling and token masking is particularly computationally efficient in contrast to the exhaustive generation and processing of tokens from the entire spectrum of information. However, randomly choosing datasets and spectral bands at each training iteration results in highly heterogeneous batches, potentially destabilizing the optimization process. To address this, the FoMo-Net pre-training process exploits gradient accumulation (see Fig. 4). Letting be a neural network parameterized by , and be a loss function, the gradient is accumulated during forward passes according to , and it is this sum that is backpropagated through .
FoMo-Net architecture: In this work, our encoder is a vision transformer (ViT) [11] with 12 layers, each with 16 attention heads, and the decoder is a ViT with 8 layers, each with 16 attention heads. We use a spatial patch size of pixels. Compared to existing foundation models, FoMo-Net is quite shallow, intended as a first step towards a generic forest-monitoring foundation model, and is designed to be lightweight since most stakeholders implementing forest-monitoring algorithms in practice have very limited computational resources. In particular, FoMo-Net1 contains 101M parameters while FoMo-Netm 110M. Both FoMo-Net versions are pre-trained for 80 epochs.
Challenges: First, FoMo-Net should model interactions between a wide range of spectral bands that could potentially be utilized in a downstream task, contrary to standard foundation models that are tailored for specific sensors. Secondly, modeling data with varying GSDs require features and patterns at different scales whereas some of them are a) underrepresented and b) lack data points with wide spatial coverage. Finally, although FoMo-Net employs a data-driven strategy with limited modeling capacity since scaling up these types of architectures significantly increases the demand for computational resources.
Experimental baseline: We test the FoMo-Net pre-training framework using satellite and aerial datasets from both the FoMo-Bench and open-access satellite archives (see Table 2). We evaluate FoMo-Net on FoMo-Bench by freezing all encoder parameters and fine-tuning a task-specific head for each downstream task. We also provide an example of the potential of fully finetuning our best performing FoMo-Net on BigEarthNet-MM. Results are presented in Table 4, Table 3, and Table 5.
6 Experiments
In this section, we evaluate a range of baselines across FoMo-Bench, addressing three distinct types of computer vision tasks: classification, semantic segmentation, and object detection.
Evaluation protocols: For each task, we train a set of standard computer vision models appropriate to the task. Importantly, each of these baselines is trained in a supervised fashion on the data for that single task. This is because the heterogeneity of the tasks in both input and output data prevents the naive use of a single architecture trained on all data. (It is this challenge that we address in Section 5 with the introduction of the foundation model baseline FoMo-Net.) In the case of multi-label classification and object detection, we also evaluate a simple approach using supervised learning on multiple datasets. PolyNet denotes a ResNet-50 model pretrained on several datasets simultaneously and then finetuned and evaluated either in-domain (denoted PolyNet), i.e. on a dataset included in pretraining, or out-of-domain (denoted PolyNet), i.e. on a dataset not included in pretraining.
During the finetuning process of in-domain, out-of-domain and FoMo-Net backbones we freeze the encoder. For classiication we use a linear evaluation protocol, whereas for object detection and semantic segmentation we train a task-specific head to learn the downstream task for each dataset independently. The FoMo-Net backbone remains the same for all experiments presented in Tables 4, 3 and 5. For all experiments, we present the average performance along with the standard deviation calculated across three independent runs using different random seeds. Detailed training settings for all experiments will be provided in Appendix B.
Semantic segmentation of grid-based data: FoMo-Bench includes seven datasets suited to semantic segmentation: Waititu, ForestNet, Spekboom FLAIR #1, FLAIR #2, Woody, and FiveBillionPixels. Our baseline models here are UNet [56], DeepLabv3 [7] and UPerNet [69]. We use a Swin Transformer as a backbone for UPerNet and a ResNet50 for the rest.
| Dataset | Modalities | #C | Method | #P (Mil.) | F1 micro | mIoU |
| Waititu [33] | Aerial RGB | 3 | UNet | 32.52 | 77.96 () | 64.0 () |
| DeepLabv3 | 26.67 | 80.81 () | 67.90 () | |||
| UPerNet | 121,30 | 75.36 () | 60.56 () | |||
[…]
| FoMo-Netm | 15.70 | 71.52 () | 55.67 () | |||
| FLAIR #1 [20, 19] | Aerial multisp. | 19 | UNet | 32.52 | 72.80 () | 57.24 () |
| DeepLabv3 | 26.68 | 74.01 () | 58.75 () | |||
| UPerNet | 121.30 | 63.89 () | 46.94 () | |||
[…]
| FoMo-Netm | 25.142 | 36.6 () | 22.43 () | |||
| FLAIR #2 [18, 19] | Aerial multisp. | 19 | UNet | 32.56 | 72.99 () | 57.50 () |
| DeepLabv3 | 26.7 | 83.97 () | 72.37 () | |||
| UPerNet | 121.33 | 76.93 () | 62.52 () | |||
[…]
| FoMo-Netm | 72.32 | 25.07 () | 14.33 () | |||
| Spekboom [17] | Aerial RGB | 2 | UNet | 32.52 | 97.70 () | 95.50 () |
| DeepLabv3 | 26.67 | 97.57 () | 95.27 () | |||
| UPerNet | 121.30 | 97.81 () | 95.72 () | |||
[…]
| FoMo-Netm | 15.70 | 85.84 () | 75.19 () | |||
| FiveBillionPixels [64] | Optical | 22 | UNet | 32,52 | 74.40 () | 59.24 () |
| DeepLabv3 | 26.68 | 73.75 () | 58.42 () | |||
| UPerNet | 121.31 | 67.43 () | 51.02 () | |||
[…]
| FoMo-Netm | 20.42 | 42.45 () | 26.94 () | |||
| ForestNet [29] | Multispec. | 5 | UNet | 32.52 | 95.26 () | 90.96 () |
| DeepLabv3 | 26.68 | 95.27 () | 90.97 () | |||
| UPerNet | 121.31 | 95.23 () | 90.90 () | |||
[…]
| FoMo-Netm | 15.70 | 95.25 () | 90.93 () | |||
| Woody [35] | Aerial RGB | 4 | UNet | 32.52 | 93.37 () | 87.57 () |
| DeepLabv3 | 26.67 | 93.30 () | 87.44 () | |||
| UPerNet | 121.30 | 93.87 () | 88.46 () | |||
[…]
| FoMo-Netm | 15.70 | 87.91 () | 78.43 () |
For UNet and DeepLabv3 architectures, we use models pretrained on ImageNet. The UPerNet is based on a Swin-Base backbone from [9]. We report the mean intersection over union (mIoU) and the F1-Score at micro level for all datasets. Results are presented in Table 3.
Multi-label classification: Due to the large ground sampling distance of sensors in remote sensing, classification tasks are usually multi-label. Given a sample image, the goal is to predict all classes that are present in that image. The multi-label classification setting is thus used for BigEarthNet-MM, Sen12MS, RapidAI4EO, TreeSat AI and TalloS for either LULC or tree species recognition. Our baselines for the classification tasks include ResNet50 [26], ViT [11] and ConvNext [40] models. We initialize all models with weights pretrained on ImageNet. We pretrain PolyNet on four classification datasets within FoMo-Bench (RapidAI4EO, TalloS, Sen12MS and BigEarthNet), with their common bands (B2, B3, B4, B5, B6, B7, B8, B8a and B11 from Sentinel-2). In-domain evaluation is performed on these datasets, while TreeSatAI is excluded from pre-training to assess the encoder’s out-of-domain generalization abilities. We report micro and macro F1 scores for all datasets in Tab. 4.
Object detection The object detection task is useful in forest monitoring to count and precisely locate trees, focusing in some cases on particular species. As it requires distinguish the canopies of different trees, aerial RGB data are particularly well suited for this task. We use the NeonTree and ReforesTree datasets to test binary and multi-class object detection respectively, testing the Faster R-CNN [54], RetinaNet [39] and YOLOS [13] models (see Table 5). We measure performance by the mean intersection over union (mIoU), the mean average precision with an threshold () and the mean average precision with a variable threshold of ().
@@ 6 Experiments
| BigEarthNet [59] | Multispec. SAR | 19 | ResNet50 | 23.58 | 77.18 () | 70.32 () |
| ViT | 87.97 | 64.42 () | 47.64 () | |||
@@
| FoMo-Netm | 0.019 | 29.15 () | 14.13 () | |||
[…]
| Sen12MS [57] | Multispec. SAR | 17 | ResNet50 | 23.57 | 51.34 () | 32.62 () |
@@
| FoMo-Netm | 0.017 | 11.02 () | 14.29 () | |||
| TalloS (ours) | Multispec. SAR DEM | 1160 | ResNet50 | 26.34 | 18.36 () | 0.68 ( |
| ViT | 89.20 | 8.52 () | 0.24 () | |||
| ConvNext | 88.98 | 8.62 () | 0.25 () | |||
| PolyNet | 2.79 | 5.81 () | 0.19 () | |||
@@
| FoMo-Netm | 1.39 | 7.29 () | 0.20 () | |||
| RapidAI4EO [2] | Optical Multispec. | 24 | ResNet50 | 23.60 | 62.84 () | 46.33 () |
| ViT | 88.37 | 31.61 () | 7.5 () | |||
@@
| FoMo-Netm | 0.026 | 12.12 () | 2.58 () | |||
| TreeSatAI [1] | Multispec. | 15 | ResNet50 | 23.56 | 53.01 () | 30.19 () |
@@
| FoMo-Netm | 0.015 | 0.04 () | 0.019 () |

### v3

@@ 4 The TalloS Dataset
As shown in Fig. 2, the spatial coverage of pre-existing datasets in FoMo-Bench is limited. We improve the diversity and spatial coverage of FoMo-Bench by designing a new dataset, TalloS. This dataset couples manual tree inventories with multispectral and SAR timeseries imagery for a fine grained multi-label classification challenge. As described in the previous section, various forest monitoring tasks have relied on high-resolution UAV data, which, despite offering detailed insights, suffer from limited spatial coverage due to human and hardware constraints (Galuszynski et al. 2022; Reiersen et al. 2022; Kattenborn et al. 2019, 2020). TalloS is a large-scale multi-modal dataset of satellite-based imagery for tree species multi-label classification. It is based on the global Tallo database (Jucker et al. 2022), containing almost manual georeferenced recordings aggregating information from forest inventories including the tree species as well as tree height, stem diameter, and crown radius.
| Dataset | Sensors | Spatial res. | Sampling Weight |
@@
| RapidAI4EO | Planet, S-2 | Medium, Low | 0.2 |
| TalloS | S-1, S-2, DEM | Low | 0.2 |
@@
Due to the modest GSD of publicly accessible satellites (), we design TalloS to concentrate on areas containing a minimum of 50 trees, as regions with lower tree populations lack distinctive signatures in satellite imagery. This choice balances diversity and strong signal for model prediction. There are k locations including 50 or more trees, but only k with 100 or more, reducing the dataset size by 25.1% and the number of genera by more than 11%. TalloS data represent a time series from 2016 to 2023 with a temporal frequency of 15 days. Each timestep consists of a multispectral image from Sentinel-2, a SAR image from Sentinel-1, a digital elevation model (DEM) from Copernicus GLO-30 at 10m per pixel resolution, along with meteorological information from ERA-5. Sentinel-2 data is selected with cloud coverage below 30%, and Sentinel-1 captures are aligned as closely as possible to Sentinel-2 acquisitions. Spectral bands with lower GSD are resampled to 10m per pixel for consistency with conventional models. Each RS sample of TalloS is mapped to ground-validated and fine grained measurements of trees in its given area, including tree species, genus and family as well as their specific traits. For the purpose of this study, we focus on the genus level labels, as distinguishing between species may require higher resolution imagery. Throughout this paper, unless explicitly stated otherwise, we refer to the various genus categories of TalloS as its classes. TalloS represents a fine-grained classification problem involving 1,000 classes with a heavily skewed data distribution. The task is framed as multi-label classification, with the goal of predicting every tree genus that is depicted in the image timeseries. In Fig. 2(b), we illustrate the wide spatial distribution of TalloS as well as the proposed training, validation and test splits. Our goal is to maximize the geographic spatial coverage of each split while ensuring that (a) regions used for training and testing do not overlap, (b) most classes present in the validation and test sets are also included in the training set. Accommodating both conditions proved challenging due to the limited geographic distribution of many tree genera. We present TalloS samples along with its class distribution in Appendix A.
5 FoMo-Net
Building on FoMo-Bench and inspired by the necessity for a generalized foundation model, we introduce FoMo-Net. FoMo-Net is a pretraining paradigm for learning to process all of the most common modalities in the RS domain with a single, sensor-agnostic foundation model. In this work we treat each spectral band of each sensor as a separate modality. Our approach is driven by three core considerations. First, it should maximize flexibility to process various input settings, without relying on specific sets of sensory modalities. Second, it should have the capacity to process information and generate meaningful representations for the whole globe. Third, it should be applicable to a wide range of downstream tasks.
@@ 5 FoMo-Net
Pre-training data: FoMo-Net pre-training scheme contains rich multi-sensor information, both paired and unpaired, from most parts of the world. In particular, we use the RapidAI4EO dataset as a source of Planet and Sentinel-2 data; TalloS containing DEM, Sentinel-1, and Sentinel-2; SSL4EO-Landsat (Stewart et al. 2024) to acquire global information from Landsat 8 and 9; a combination of all available UAV datasets in FoMo-Bench; and the FiveBillionPixel dataset providing a unique source of Gaofen-2 high resolution satellite imagery. The datasets used in the FoMo-Net pre-training framework are detailed in Tab. 1. The proposed framework is able to process any combination of the 36 most common modalities in RS, stemming from Sentinel-1, Sentinel-2, Landsat 8-9, Planet, Gaofen-2 and UAV sensors with ground sampling distance spanning from a few centimeters to per pixel. In this work, a distinct modality is defined as any band with a unique combination of wavelength and spatial resolution. Given its importance in Earth observation, digital elevation models (DEM) are also included in our definition as a specific band. The FoMo-Net pre-training pipeline is illustrated in Fig. 3 and will be detailed in the following sections. Additionally, the pseudocode of the pre-training pipeline is summarized in Appendix B. Point-cloud data are excluded from the pretraining set and initial FoMo-Net design due to the need for specialized processing, which is deferred to future work (Sec. 7).
@@
Approach to variable spectral bands: Let be a set of datasets, and be a set of spectral bands, where where and represent respectively the height and width dimensions. At each iteration, a training batch contains a variable number of spectral bands, each sampled with respective probabilities . In our experiments, we set proportional to the frequency with which the -th band occurs in (see Tab. 1).
@@
Each band is tokenized into patches of size according to the transformation , so that . The tokenized input is then embedded using a linear transformation learnt during the optimization process as . We considered two projection setups: The first (noted FoMo-Net1) uses a single linear projection for all spectral bands, while the second (denoted FoMo-Netm) projects each spectral band with its own linear projection. Our initial experiments suggest that FoMo-Net1 yields significantly better results. The input to our model, then, is provided as:
@@
Spectral band MAE: The next step of the procedure follows a masked autoencoding (MAE) framework (He et al. 2022) that is shared by various high-performing foundation models for RS (see Sec. 2). However, our method differs from a typical MAE in two key aspects (see Fig. 3).
@@
First, we introduce a random spectral band selection process by sampling in both and , representing the datasets and spectral bands, respectively. The number of selected spectral bands is randomly defined at each iteration. This approach promotes modeling interactions between a wide range of band combinations, creating a highly flexible encoder able to process any set of spectral bands. Next, we incorporate both spectral and positional embeddings, allowing the model to capture information regarding the spatial position of the patch and its spectral band, as detailed in Eq. 1. Both types of embeddings are trainable and randomly initialized. Given the randomness in the input space, the spectral embedding is of utmost importance as no other information regarding the nature of the input is provided. Finally, as each band is tokenized independently the transformer is able to apply the attention mechanism to both spectral and spatial dimensions.
@@
Batch gradient accumulation: The combination of the proposed modality sampling and token masking is particularly computationally efficient in contrast to the exhaustive generation and processing of tokens from the entire spectrum of information. However, randomly choosing datasets and spectral bands at each training iteration results in highly heterogeneous batches, potentially destabilizing the optimization process. To address this, the FoMo-Net pre-training algorithm exploits gradient accumulation (see Fig. 3). Letting be a neural network parameterized by , and be a loss function, the gradient is accumulated during forward passes according to , where is the -th batch of input data (see Eq. 1), and it is this sum that is backpropagated through .
FoMo-Net backbone: In this work we apply FoMo-Net on a ViT encoder with 12 layers, and 12 attention heads. We limit the maximum number of sampled spectral bands at the pre-training stage to 4 and the pixel resolution to for efficiency and train for 300 epochs.
Modelling challenges: FoMo-Net must account for interactions across a wide range of spectral bands potentially useful for downstream tasks, unlike standard foundation models tailored to specific sensors. Additionally, varying GSD requires capturing features and patterns at multiple scales, some of which are a) underrepresented and b) lack sufficient data points with wide spatial coverage. Finally, given the diversity of information in the input and the large-scale pretraining dataset, scaling up the backbone could further improve performance. However, this would require significantly more computational resources.
@@
| BigEarthNet (Sumbul et al. 2021) | Multispec. SAR | 19 | ResNet50 | 23.58 | 77.18 () | 70.32 () |
| ViT | 87.97 | 78.09 () | 72.29 () | |||
@@
| FoMo-Netm | 73.57 | 66.11 () | 50.0 () | |||
| Sen12MS (Schmitt et al. 2019) | Multispec. SAR | 17 | ResNet50 | 23.57 | 51.34 () | 32.62 () |
@@
| FoMo-Netm | 73.57 | 38.07 () | 19.44 () | |||
| TalloS (ours) | Multispec. SAR DEM | 1160 | ResNet50 | 26.34 | 27.31 () | 1.35 () |
| ViT | 89.20 | 15.23 () | 0.37() | |||
| ConvNext | 88.98 | 9.08 () | 0.43 () | |||
| PolyNet | 2.79 | 4.57 () | 0.13 () | |||
@@
| FoMo-Netm | 74.45 | 14.25 () | 0.27 () | |||
| RapidAI4EO (Bhugra et al. 2022) | Optical Multispec. | 24 | ResNet50 | 23.60 | 62.84 () | 46.33 () |
| ViT | 88.37 | 51.04 () | 25.59 () | |||
@@
| FoMo-Netm | 73.58 | 49.12 () | 24.63 () | |||
| TreeSatAI (Ahlswede et al. 2023) | Multispec. | 15 | ResNet50 | 23.56 | 53.01 () | 30.19 () |
@@
| FoMo-Netm | 73.57 | 27.96 () | 12.08 () |
6 Experiments
We train a set of standard computer vision models tailored to each specific task. Each baseline is trained in a supervised fashion on the data for that single task due to the heterogeneity of the tasks in both input and output. In the case of multi-label classification and object detection, we also evaluate a simple approach using supervised learning on multiple datasets i.e. PolyNet. PolyNet denotes a ResNet-50 model pretrained on several datasets simultaneously. Since PolyNet has been trained in a supervised fashion on FoMo-Bench tasks, we evaluate it via linear probing either in-domain (denoted PolyNet), i.e. on a dataset included in pre-training, or out-of-domain (denoted PolyNet), i.e. on a dataset not included in pre-training. For object detection and semantic segmentation we train a task-specific head to learn the downstream task for each dataset independently. For all experiments, we present the average performance along with the standard deviation over three independent runs using different random seeds. Detailed training settings and qualitative results are provided in Appendices C and D respectively.
Semantic segmentation of grid-based data: FoMo-Bench includes seven datasets suited to semantic segmentation. Our baseline models are UNet (Ronneberger, Fischer, and Brox 2015), DeepLabv3 (Chen et al. 2017) and UPerNet (Xiao et al. 2018).
| Dataset | Modalities | #C | Method | #P (Mil.) | F1 micro | mIoU |
| Waititu (Kattenborn et al. 2020) | Aerial RGB | 3 | UNet | 32.52 | 77.96 () | 64.0 () |
| DeepLabv3 | 26.67 | 80.81 () | 67.90 () | |||
| UPerNet | 121,30 | 75.36 () | 60.56 () | |||
[…]
| FoMo-Netm | 86.49 | 71.52 () | 55.66 () | |||
| FLAIR #1 (Garioud et al. 2022, 2024) | Aerial multisp. | 19 | UNet | 32.52 | 72.80 () | 57.24 () |
| DeepLabv3 | 26.68 | 74.01 () | 58.75 () | |||
| UPerNet | 121.30 | 63.89 () | 46.94 () | |||
[…]
| FoMo-Netm | 93.57 | 60.22 () | 43.08 () | |||
| FLAIR #2 (Garioud et al. 2023, 2024) | Aerial multisp. | 19 | UNet | 32.56 | 72.99 () | 57.50 () |
| DeepLabv3 | 26.7 | 83.97 () | 72.37 () | |||
| UPerNet | 121.33 | 76.93 () | 62.52 () | |||
[…]
| FoMo-Netm | 128.96 | 57.95 () | 40.80 () | |||
| Spekboom (Galuszynski et al. 2022) | Aerial RGB | 2 | UNet | 32.52 | 97.70 () | 95.50 () |
| DeepLabv3 | 26.67 | 97.57 () | 95.27 () | |||
| UPerNet | 121.30 | 97.81 () | 95.72 () | |||
[…]
| FoMo-Netm | 86.49 | 91.83 () | 70.97 () | |||
| FiveBillionPixels (Tong, Xia, and Zhu 2023) | Optical | 22 | UNet | 32,52 | 74.40 () | 59.24 () |
| DeepLabv3 | 26.68 | 73.75 () | 58.42 () | |||
| UPerNet | 121.31 | 67.43 () | 51.02 () | |||
[…]
| FoMo-Netm | 90.03 | 67.39 () | 50.82 () | |||
| ForestNet (Irvin et al. 2020) | Multispec. | 5 | UNet | 32.52 | 95.26 () | 90.96 () |
| DeepLabv3 | 26.68 | 95.27 () | 90.97 () | |||
| UPerNet | 121.31 | 95.23 () | 90.90 () | |||
[…]
| FoMo-Netm | 86.49 | 95.24 () | 90.92 () | |||
| Woody (Kattenborn et al. 2019) | Aerial RGB | 4 | UNet | 32.52 | 93.37 () | 87.57 () |
| DeepLabv3 | 26.67 | 93.30 () | 87.44 () | |||
| UPerNet | 121.30 | 93.87 () | 88.46 () | |||
[…]
| FoMo-Netm | 86.49 | 84.33 () | 72.93 () |
For UNet and DeepLabv3 architectures, we use a ResNet50 (He et al. 2016) backbone pretrained on ImageNet. The UPerNet is based on a Swin-Base backbone from (MMSegmentation Contributors 2020). We assess FoMo-Net by extracting the encoded tokens, upsampling and feeding them to a small decoder similar to FloodViT (Bountos et al. 2024). We report the mean intersection over union (mIoU) and the F1-Score at micro level for all datasets. Results are presented in Tab. 3.
Classification: Due to the large GSD of sensors in RS, classification tasks are usually multi-label. The multi-label classification setting is thus used for BigEarthNet-MM, Sen12MS, RapidAI4EO, TreeSat AI and TalloS for either LULC or tree species recognition. Our baselines are ResNet50, ViT and ConvNext (Liu et al. 2022) models. We initialize all models with weights pretrained on ImageNet. We pretrain PolyNet on RapidAI4EO, TalloS, Sen12MS and BigEarthNet, with their common bands. In-domain evaluation is performed on these datasets, while TreeSatAI is used for out-of-domain evaluation.We report micro and macro F1 scores for all datasets in Tab. 2.
Object detection The object detection task is useful in forest monitoring to count and precisely locate trees, focusing in some cases on particular species. We use the NeonTree and ReforesTree datasets to test binary and multi-class object detection respectively, testing the Faster R-CNN (Ren et al. 2016), RetinaNet (Lin et al. 2020) and YOLOS (Fang et al. 2021) models (see Tab. 4). We measure performance by the mIoU, the mean average precision with an threshold () and the mean average precision with a variable threshold of ().